OpenStudy (rayman5766):

what is 40 approxmantly to the nears tenth A 5.8 B 6.3 C 6.0 D 6.7

OpenStudy (rayman5766):

please help me i have been on one question for a hole hour

OpenStudy (welshfella):

I dont understand the question. r u sure its 40?

OpenStudy (rayman5766):

oh right its square root 40

OpenStudy (rayman5766):

sorry

OpenStudy (welshfella):

ahh OK

OpenStudy (welshfella):

well the correct choice can be found if you square them to see which comes nearest to 40 Use ur calculator.

OpenStudy (rayman5766):

i know that but which number will i use 8 x 5

OpenStudy (rayman5766):

well wait thats 40

OpenStudy (welshfella):

6^2 = 36 so you eliminate this and 5.8

OpenStudy (welshfella):

are you studying methods of finding square roots at the moment?

OpenStudy (rayman5766):

dont you need to number like this 6<5.8<8

OpenStudy (rayman5766):

are something similar to this

OpenStudy (rayman5766):

also yes yes i am

OpenStudy (welshfella):

OH Theres only one I know - not sure if i can remember it! With regard to this problem the root is between 6 and 7 because 40 is between 36 and 49. now you use proportion to estimate it more accurately Is that what you are using?

OpenStudy (rayman5766):

well 36 is obviously closer

OpenStudy (welshfella):

49 - 36 = 13 and 40 - 36 = 4 so estimate of the root is 6 + 4/13

OpenStudy (rayman5766):

82/13

OpenStudy (rayman5766):

which 6.03

OpenStudy (welshfella):

yes now convert that to decimal form

OpenStudy (rayman5766):

well 6.30

OpenStudy (rayman5766):

so its 6.3

OpenStudy (welshfella):

yes 6.3 is correct

OpenStudy (rayman5766):

thanks bud i understand this alot better now

OpenStudy (rayman5766):

wait

OpenStudy (rayman5766):

before you go i have to get near it not on it right

OpenStudy (welshfella):

yes - this method only gives approximate answers.

OpenStudy (rayman5766):

oh right i forgot about that word

OpenStudy (rayman5766):

thanks welshfella

OpenStudy (welshfella):

its pretty god - if you find sqrt 40 on calculator it gives 6.325

OpenStudy (welshfella):

yw

OpenStudy (rayman5766):

was that a cat noise yw